The revenue for a sandwich shop is directly proportional to its advertising budget. when the owner spent $2000 a month on advertising, the revenue was $120,000. If the revenue is now $180,000, how much is the owner spending on advertising?
Answer by bucky(2189)

Let's use a little logical reasoning.
If the owner got $120,000 in revenue by spending $2000 for ads, then we set up the ratio:
120,000/2000
but this ratio will be constant throughout the range. So we can ask if 180,000 is the sales, what amount will have to be spent on ads to get the same ratio?
We can set up a proportion:
120,000/2000 = 180,000/x
Notice that the sales are the numerators of both sides and the advertising costs are the denominators of both sides.
Multiply both sides by x to get the x into a numerator:
120,000x/2000 = 180,000
Multiply both sides by 2000 to eliminate the denominator on the left side:
120,000x = 180,000 * 2000
Divide both sides by 120,000 to solve for x which is the advertising dollars for $180,000 in sales:
x = (180,000 * 2000)/120,000
Do the multiplication and division on the right side to find that the advertising expenses associated with $180,000 in sales is $3,000.
Hope this helps.